A perfectly mixed aeration pond with no recycle (return line) serves as the biological reactor for a small community. The pond receives 30 m3/d of influent with a BOD5 of 350 mg/L that must be reduced to 20 mg/L before discharge. It has been found that the kinetics constant for the system are Ks = 100 mg/L BOD5, kd = 0.10 d-1, um = 1.6 d-1, and Y is 0.60 mg BOD5.

 

 A)What must the hydraulic detention time be in the aeration pond?

 B) What mass of microbes will be produced in the pond each day?


1
Expert's answer
2021-07-07T06:21:02-0400

Given data:

Final lelvel of BOD is "20 mg\/l"

"K_s = 100 mg BOD\/L,"

"k_d = 0.10 day^{-1} ,"

"\u03bcm = 1.6 day^{-1} ,"

"Y = 0.60 mg SS\/mg BOD"

a) we know that "C_{out}=\frac{C_{in}}{1+kt}"

"20=\frac{350}{1+0.1*t} \implies t = 165"

solving for t

"t = 165 days"

b) mass of microbes 

"Q(mlD)AC(mg\/l)"

"3*10^{-3}(350-20)=9.9kg\/day"